# Framing in the Presence of Supporting Data: A Case Study in U.S. Economic News

### v2

@@ 1 Introduction
In this paper, we make the observation that there are a select number of newsworthy topics where objective measures do exist. For example, there is a finite set of economic indicators that experts use to define the state of the economy, such as changes in nonfarm payroll and gross domestic product. Another example is crime, which has well-established data points maintained by the FBI and local jurisdictions. Building a system capable of tracking a large number of publications, identifying which indicators are reported and how they are being reported has enormous research potential. Such a system would enable us create three views of potential bias: how a publication is covering a topic over time, how their coverage differs from other publications, and how it compares to the accepted ground truth of experts in the field.
@@
To address this problem, we decompose economic frames into a set of interdependent tasks (See Fig. 1). At the article level, we learn to identify the relevance of the article to the economy, the type of economic information it covers (e.g., macro economic, industry-specific, firm-specific), and two measures of general economic sentiment: the state of the economy (e.g., good, fair, poor), and the direction in which the economy is heading (e.g., better, same, worse). Then, for each numerical quantity reported in the article, we learn to predict whether it corresponds to an economic indicator (e.g., market numbers) and its reported polarity (e.g., “The S&P 500 fell 3 percent” has a negative polarity). This decomposition captures three levels: whether the economy is being discussed, which economic indicators are highlighted, and the framing of those indicators and the economy as a whole. This way, we can go beyond analyzing topic selection by additionally measuring how a publication’s reporting differs from the objective reality at that time, how it evolves across time, and how it compares to other publications.
We make the following contributions: (1) We propose a computational framework to model framing in the presence of supporting data and apply it to news about the U.S. economy. (2) We collect a novel dataset of landing page news articles published by a set of major U.S. news outlets between 2015 and 2023, and provide high-quality annotations of our proposed frame components for a small subset of examples. (3) We propose an automated method to detect each of the proposed frame components under low-supervision settings. (4) We demonstrate that our framework can be used to track when and how often (selection), as well as how (framing) different news outlets report on specific aspects of the economy. We compare these reports to the ground truth for an example indicator: job numbers. All of our code and data will be released to the community.
@@ 2 Background and Related Work
Theory of Selection and Framing
National news publications have a massive choice-set of topics they could cover any given day: it may feel like the publications are reacting to a defined set of events that recently happened but, outside of a few major events, they choose what the news is on any given day. In other words, the news in mainstream publications does not present an objective view of current affairs. Rather, it is influenced by the selection made by media professionals McQuail (1992). Selection is driven by what researchers have called news values Galtung and Ruge (1965); Harcup T and O’Neill D (2001, 2017), which determine whether a story is newsworthy. While there is variation in the taxonomies that have been proposed to determine newsworthiness, the most recent ones include factors such as: the power of elite (stories concerning powerful entities), bad and good news (stories with particular emotional overtones), and magnitude (stories that are perceived as significant for a wide audience) Harcup T and O’Neill D (2017).
More interestingly, news values vary from publication to publication. A study of the run-up to the 2022 election showed that the New York Times and Washington Post averaged just two overlapping topics each morning in their printed editions, out of about six articles Rothschild et al. (2023).
Conditional on selecting a topic, there is additional leeway on how to cover it, as there is a wide-range of facts and opinions that could spin the takeaway of the readers. This phenomenon, referred to as framing, has been widely studied in communication studies. While there is not an agreed upon definition of framing, recent surveys Vallejo et al. (2023) have identified the following prevalent definitions: equivalence framing (presenting the same exact information in different ways) Cacciatore et al. (2015), emphasis framing (highlighting specific aspects of an event to promote a particular interpretation) Entman (2007), and story framing (leveraging established narratives to convey information) Hallahan (1999). When it comes to the economy, a publication could present a rosy or terrible version of the economy, without lying, just by picking the right economic indicator with the right perspective. For example, there are months where non-farm payroll is up more than the average but it is below the market expectation: either perspective is technically true. But, with a relatively finite set of indicators, and an established way for economists to judge them, there is still a clear ground truth to compare if any given indicator with any given perspective represents the consensus understanding of the economy at that time.
@@
Understanding the health and wealth of the U.S. economy is critical for investment and growth. Most of the modern economic indicators have been tracked for decades, codified in the by the mid-20th century. Example include: jobs numbers (e.g., nonfarm payroll), prices (e.g., consumer price index), and macro economic (e.g., gross domestic product). With the exception of market prices (e.g., S&P500), which are a reflection of the underlying assets, the core indicators are tracked by a few government agencies (e.g, the Bureau of Economic Analysis for GDP, spending, income, trade, and the Bureau of Labor Statistics for unemployment, prices, productivity), which conduct regular, large scale polling.
These economic indicators inform the public on the health of the economy. The jobs numbers are front page news when they are released at 8:30 AM ET on the first Friday of each month, the state of the markets is constantly in the news, and indicators such as the consumer price index pop up whenever the media finds them most important, interesting, and engaging. Most people have limited views of the economy outside of their close social circle, so these values are the key to their understanding of the health of the economy as a whole. And, traditionally, the health of the economy has tracked very closely with the approval of incumbent political leaders Hummel and Rothschild (2014).
@@
This paper builds off an extensive literature that has not just identified the quantity of economic coverage, but framing of that coverage. Most of this work is geared towards matching the sentiment of articles about the economy to both consumer sentiment and economic indicators, tracking both historical value and predicting future ones Hopkins et al. (2017); Ardia et al. (2019); Shapiro et al. (2022); Seki et al. (2022); Bybee (2023); van Binsbergen et al. (2024). While the methods continue to evolve from simple keywords to BERT or LLM-based predictions, they are focused on article-level sentiment, aggregated to states or the U.S. They also study a mix of questions about economic impact that benefit from a long time-series of sentiment, granular in both time and geography, regarding questions as diverse as the impact of bubbles on industries (Bybee, 2023), economic shocks to economic decisions Shapiro et al. (2022), to the relationship between news and perception of the economy Hopkins et al. (2017). Our paper extends this literature by not just capturing sentiment on economic articles, but what economic indicators and economic indicator level sentiment is driving it. This allows us to dissect what editorial choices are driving perceived sentiment around the economy
@@
With the proliferation of NLP techniques, scholars have increasingly adopted computational approaches to study framing in diverse forms of media. By automating parts of the process, studies have been able to scale to very large media repositories. In most cases, researchers adopt unsupervised techniques such as topic modeling to identify latent themes DiMaggio et al. (2013); Nguyen et al. (2015); Gilardi et al. (2021). However, topic models limited in their ability to capture framing nuances. By defining topics as simple distributions over words, they lack the semantic and discursive conceptualization needed to answer how issues and events are being presented Ali and Hassan (2022).
On the other hand, there is a body of work leveraging supervised learning Johnson et al. (2017); Khanehzar et al. (2019); Kwak et al. (2020); Huguet Cabot et al. (2020); Mendelsohn et al. (2021) and lexicon expansion techniques Field et al. (2018); Roy and Goldwasser (2020) to analyze framing. For this to be feasible, authors need to define a concrete taxonomy of which frames are relevant and how they are represented in data. Most of these approaches rely on the taxonomy proposed by  Boydstun et al. (2014). These dimensions correspond to broad themes such as “economic”, and “public opinion”. While these themes have shown to be useful to model a select set of contentious political issues like abortion and immigration, they are too broad to capture how the economy and other data-driven topics are being presented in the media.
@@ 3 Data
To identify articles that discuss the economy, we curated a lexicon of economic terms sourced from the Bureau of Labor Statistics and the FRED database operated by the Federal Reserve bank of St. Louis.***https://www.bls.gov/data/https://fred.stlouisfed.org/ These sources provide a fairly comprehensive list of common metrics of economic activity. We consider an article to be relevant to the economy if it contains at least three sentences mentioning an economic term from our lexicon. This resulted in a total of 199,066 articles (Tab. 1). App. A.1 includes additional details about the data collection process and the lexicon of economic terms.
@@
We employed a group of six annotators to label our dataset following the guidelines outlined in App. A.3. Our annotators comprised senior, postdoctoral and pre-doctoral researchers. Their fields of study include: economics, computational social science, and communications. They received training from the authors of the paper and several rounds of calibration were performed before annotation began. Articles were chosen such that there would be some overlap between coders (in order to check quality), as well as a decent breadth of coverage. To promote topical breadth, we followed the sampling procedure introduced in Pacheco et al. (2022b, 2023).
For each article in their batch, annotators selected: (1) what type of economic information was the most prominent (macro, government, industry-specific, business-specific or personal), (2) the framing of the general economic conditions (good, fair, bad), and (3) the framing of the direction that the economy is heading (better, same, worse). These questions were adapted from the Gallup economic index https://news.gallup.com/poll/1609/consumer-views-economy.aspx. Then, for every valid numerical value reported in the article, annotators selected: (1) what type of economic data it reports (macro, government, industry-specific, business-specific or personal), (2) which indicator was reported (e.g., jobs, prices. See App. A.3 for the full list of options), and (3) the reported polarity of each quantity (positive, negative, neutral). We report resulting statistics in Table 1, and inter-annotator agreement in Table 2. To calculate agreement, we only consider examples that were annotated by two or more coders. In App. A.4 we include statistics for the number of coders for each of the label categories. Lastly, we show the resulting annotation distribution for all frame components in Fig. 2.
@@
We find that inter-annotator agreement is generally good for quantity-level information (Krippenforff’s $\alpha$ 0.56-0.83). This reflects our intuition that capturing framing at the level of supporting data points may be a better alternative to relying on general topical markers. On the other hand, article-level judgements which aim to capture main “takeaways” are considerably harder for annotators, exhibiting moderate agreement (Krippenforff’s $\alpha$ 0.42-0.48). This is expected and in line with other high-level framing tasks Card et al. (2015); Roy et al. (2021); Mendelsohn et al. (2021). To further characterize the lower agreement values, we include confusion matrices and some examples illustrating ambiguous cases in App A.5. We find that most of the confusion lies on the following article-level judgements: (1) identifying the main type of the article, particularly deciding between macro-economic and government-specific types when both types of data points are present in the article, (2) characterizing articles that are positive towards the economy, as they tend to be more subtle than the clearly negative ones, and (3) identifying articles that suggest that the economy is doing neither better nor worse, but staying the same, which is an inherently ambiguous category. We have included a similar confusion analysis for quantitative-level judgements in App A.5.
Figure 2: Distribution of annotation labels
@@ 4 Model
To model framing, we take advantage of the relations between the values of the frame components for a given article and implement a relational model using Probabilistic Soft Logic (PSL) Bach et al. (2017). Dependencies in PSL are expressed using weighted logical rules of the form: $w_{r}:P_{1}\wedge...\wedge P_{n-1}\rightarrow P_{n}$, where $w_{r}$ indicates the importance of the rule in the model, and can be learned from data. Predicates $P_{i}$ correspond to decisions and observations. Rrules are then compiled into a Hinge-Loss Markov random field and weights are learned using maximum likelihood estimation. We consider the following rules:
Priors We explicitly model prior probabilities for all article-level (type, economic conditions and economic direction) and quantity-level (type, macro-indicator, and sentiment) components. These priors are derived from supervised classifiers trained on our labeled dataset.
@@
$(r_{2})$ Consistency between the article type, economic conditions and direction We enforce that if the article type is macro, then economic conditions and economic direction must be predicted and vice-versa: $\mathtt{Type(a,macro)}\Leftrightarrow\neg\mathtt{EconConditions(q,irrelevant)}\wedge\neg\mathtt{EconRating(q,irrelevant)}$
@@
$(r_{5})$ Dependency between neighboring quantity types Our intuition is that there are common patters in the types of consecutive quantities (e.g., sequential dependencies): $\mathtt{Precedes(q1,q2)}\wedge\mathtt{Type(q1,t1)}\Rightarrow\mathtt{Type(q2,t2)}$
@@
Pre-trained language models (LMs) are one of the most successful techniques to acquire knowledge from unlabeled text data Devlin et al. (2019). Domain-adaptive pre-training (DAPT) further enriches LMs with in-domain data Gururangan et al. (2020). In both of these pre-training stages, learning is conducted by randomly masking words in a large dataset and training the the LM to predict these words. To take advantage of these strategies, we use pre-trained RoBERTA and perform DAPT with our large unlabeled dataset (all 199,066 economic articles collected). We use this LM to fine-tune the classifiers that are used as priors in PSL. This strategy has been repeatedly used to combine PSL with strong classifiers, with consistent success  Sridhar et al. (2015); Pacheco and Goldwasser (2021); Roy et al. (2021)

### v3

@@ 1 Introduction
In this paper, we make the observation that there are a select number of newsworthy topics where objective measures do exist. For example, there is a finite set of economic indicators that experts use to define the state of the economy, such as changes in non-farm payroll and gross domestic product. Another example is crime, which has well-established data points maintained by the FBI and local jurisdictions. Building a system capable of tracking a large number of publications, identifying which indicators are reported and how they are being reported has enormous research potential. Such a system would enable us to create three views of potential bias: how a publication is covering a topic over time, how their coverage differs from other publications, and how it compares to the accepted ground truth of experts in the field.
@@
To address this problem, we decompose economic frames into a set of interdependent tasks (See Fig. 1). At the article level, we identify the relevance of the article to the economy, the type of economic information it covers (e.g., macro economic, industry-specific, firm-specific), and two measures of general economic sentiment: the state of the economy (e.g., good, fair, poor), and the direction in which the economy is heading (e.g., better, same, worse). Then, for each numerical quantity reported in the article, we predict whether it corresponds to an economic indicator (e.g., market numbers) and its reported polarity (e.g., “The S&P 500 fell 3%” has a negative polarity). This decomposition captures three levels: whether the economy is being discussed, which economic indicators are highlighted, and the framing of those indicators and the economy as a whole. This way, we can go beyond analyzing topic selection by additionally measuring how a publication’s reporting differs from the objective reality at that time, how it evolves across time, and how it compares to other publications. We present a computational model that predicts each proposed frame component jointly. It utilizes the inter-dependence between the sub-tasks (e.g., polarity of quantities and polarity of the article they appear in) to improve upon results obtained by independent classifiers trained to make predictions individually with limited supervision.
We make the following contributions: (1) We propose a computational framework to model framing in the presence of supporting data and apply it to news about the U.S. economy. (2) We collect a novel dataset of landing page news articles published by a set of major U.S. news outlets between 2015 and 2023, and provide high-quality annotations of our proposed frame components for a small subset of examples. (3) We propose an automated method to detect each of the proposed frame components under low-supervision settings. (4) We demonstrate that our framework can be used to track when and how often (selection), as well as how (framing) different news outlets report on specific aspects of the economy. We compare these reports to the ground truth for two example indicators: job numbers and prices. All of our code and data has been released to the community***https://github.com/blast-cu/econ-indicators.
@@ 2 Background and Related Work
Selection and Framing
National news publications have a massive choice-set of topics they could cover on any given day. Though it may seem like the publications are reacting to a defined set of events that recently happened, outside of a few major events, they choose what the news is on any given day. In other words, the news in mainstream publications does not present an objective view of current affairs. Rather, it is influenced by the selection made by media professionals McQuail (1992). Selection is driven by what researchers have called news values Galtung and Ruge (1965); Harcup T and O’Neill D (2001, 2017), which determine whether a story is newsworthy.
While there is variation in the taxonomies that have been proposed to determine newsworthiness, the most recent ones include factors such as: the power of the elite (stories concerning powerful entities), bad and good news (stories with particular emotional overtones), and magnitude (stories that are perceived as significant for a wide audience) Harcup T and O’Neill D (2017). More interestingly, news values vary from publication to publication. A study of the run-up to the 2022 election showed that the New York Times and Washington Post averaged just two overlapping topics each morning in their printed editions out of about six articles Rothschild et al. (2023).
Conditional on selecting a topic, there is additional leeway on how to cover it, as there is a wide-range of facts and opinions that could spin the takeaway of the readers. This phenomenon, referred to as framing, has been widely studied in communication studies. While there is not an agreed upon definition of framing, recent surveys Vallejo et al. (2023) have identified the following prevalent definitions: equivalence framing (presenting the same exact information in different ways) Cacciatore et al. (2015), emphasis framing (highlighting specific aspects of an event to promote a particular interpretation) Entman (2007), and story framing (leveraging established narratives to convey information) Hallahan (1999).
When it comes to the economy, a publication could present a rosy or terrible version of the economy. They can do so without lying by simply picking the right economic indicator with the right perspective. For example, there are months where the value of the non-farm payroll is above average, but below the market expectation: either perspective is technically true. But, with a relatively finite set of indicators, and an established way for economists to judge them, there is still a clear ground truth to compare if any given indicator with any given perspective represents the consensus understanding of the economy at that time.
@@
Understanding the health and wealth of the U.S. economy is critical for investment and growth. Most modern economic indicators have been tracked for decades, codified in the by the mid-20th century. Examples include: jobs numbers (e.g., non-farm payroll), prices (e.g., consumer price index), and macro economic (e.g., gross domestic product). With the exception of market prices (e.g., S&P500), which are a reflection of the underlying assets, the core indicators are tracked by a few government agencies (e.g, the Bureau of Economic Analysis for GDP, spending, income, trade, and the Bureau of Labor Statistics for unemployment, prices, productivity), which conduct regular, large scale polling.
These economic indicators inform the public on the health of the economy. The job numbers are front page news when they are released at 8:30 AM ET on the first Friday of each month. The state of the markets is constantly covered in the news. Indicators such as the consumer price index pop up whenever the media finds them most important, interesting, and engaging. Most people have limited views of the economy outside of their close social circle, so these values are the key to their understanding of the health of the economy as a whole. And, traditionally, the health of the economy has tracked very closely with the approval of incumbent political leaders Hummel and Rothschild (2014).
@@
This paper builds off of an extensive literature that has not just identified the quantity of economic coverage, but framing of that coverage. Most of this work is geared towards matching the sentiment of articles about the economy to both consumer sentiment and economic indicators, tracking both historical value and predicting future ones Hopkins et al. (2017); Ardia et al. (2019); Shapiro et al. (2022); Seki et al. (2022); Bybee (2023); van Binsbergen et al. (2024). While the methods continue to evolve from simple keywords to BERT or LLM-based predictions, they are focused on article-level sentiment, aggregated to states or the U.S. They also study a mix of questions about economic impact that benefit from a long time-series of sentiment, granular in both time and geography, regarding questions as diverse as the impact of bubbles on industries (Bybee, 2023), economic shocks to economic decisions Shapiro et al. (2022), to the relationship between news and perception of the economy Hopkins et al. (2017). Our paper extends this literature by capturing not only the sentiment of economic articles, but what economic indicators and economic indicator-level sentiment is driving it. This allows us to dissect what editorial choices are driving perceived sentiment around the economy.
@@
Scholars have increasingly adopted computational approaches to study framing, allowing studies to scale to large media repositories. In most cases, researchers adopt unsupervised techniques such as topic modeling to identify latent themes DiMaggio et al. (2013); Nguyen et al. (2015); Gilardi et al. (2021). However, topic models are limited in their ability to capture framing nuances. By defining topics as simple distributions over words, they lack the semantic and discursive conceptualization needed to answer how issues and events are being presented Ali and Hassan (2022).
There is also a body of work leveraging supervised learning Johnson et al. (2017); Khanehzar et al. (2019); Kwak et al. (2020); Huguet Cabot et al. (2020); Mendelsohn et al. (2021) and lexicon expansion techniques Field et al. (2018); Roy and Goldwasser (2020) to analyze framing. For this to be feasible, authors must define a concrete taxonomy of which frames are relevant and how they are represented in data. Most of these approaches rely on the taxonomy proposed by  Boydstun et al. (2014). These dimensions correspond to broad themes such as “economic”, and “public opinion”. While these themes have shown to be useful to model a select set of contentious political issues like abortion and immigration, they are too broad to capture how the economy and other data-driven topics are being presented in the media.
@@ 3 Data
To identify articles that discuss the economy, we curated a lexicon of economic terms sourced from the Bureau of Labor Statistics and the FRED database operated by the Federal Reserve bank of St. Louis.https://www.bls.gov/data/https://fred.stlouisfed.org/ These sources provide a fairly comprehensive list of common metrics for economic activity. We consider an article to be relevant to the economy if it contains at least three sentences mentioning an economic term from our lexicon. This resulted in a total of 199,066 articles (Tab. 1). App. A.1 includes additional details about the data collection process and the lexicon of economic terms.
@@
We employed a group of six annotators to label our dataset following the guidelines outlined in App. A.3. Our annotators comprised senior, postdoctoral and pre-doctoral researchers. Their fields of study include: economics, computational social science, and communications. They received training from the authors of the paper and several rounds of calibration were performed before annotation began. Articles were chosen such that there would be some overlap between coders (in order to check quality) and a decent breadth of coverage. We prioritized topic diversity over uniformity of news outlets and thus followed the sampling procedure introduced in  Pacheco et al. (2022b, 2023).
For each article in their batch, annotators selected: (1) what type of economic information was the most prominent (macro, government, industry-specific, business-specific or personal), (2) the framing of the general economic conditions (good, fair, bad), and (3) the framing of the direction that the economy is heading (better, same, worse). These questions were adapted from the Gallup economic index §§§https://news.gallup.com/poll/1609/consumer-views-economy.aspx. Then, for every valid numerical value reported in the article, annotators selected: (1) what type of economic data it reports (macro, government, industry-specific, business-specific or personal), (2) which indicator was reported (e.g., jobs, prices. See App. A.3 for the full list of options), and (3) the reported polarity of each quantity (positive, negative, neutral). We report resulting statistics in Table 1, and inter-annotator agreement in Tab.  2. To calculate agreement, we only consider examples that were annotated by two or more coders. In App. A.4 we include statistics for the number of coders for each of the label categories. Lastly, we show the resulting annotation distribution for all frame components in Figs. 2 and  3.
@@
We find that inter-annotator agreement is generally good for quantity-level information (Krippenforff’s $\alpha$ 0.56-0.83). This reflects our intuition that capturing framing at the level of supporting data points may be a better alternative to relying on general topical markers. On the other hand, article-level judgements which aim to capture main “takeaways” are considerably harder for annotators, exhibiting moderate agreement (Krippenforff’s $\alpha$ 0.42-0.48). This is expected and in line with other high-level framing tasks Card et al. (2015); Roy et al. (2021); Mendelsohn et al. (2021). To further characterize the lower agreement values, we include confusion matrices and some examples illustrating ambiguous cases in App. A.5. We find that most of the confusion lies on the following article-level judgements: (1) identifying the main type of the article, particularly deciding between macro-economic and government-specific types when both types of data points are present in the article, (2) characterizing articles that are positive towards the economy, as they tend to be more subtle than the clearly negative ones, and (3) identifying articles that suggest that the economy is doing neither better nor worse, but staying the same, which is an inherently ambiguous category. We have included a similar confusion analysis for quantitative-level judgements in App. A.5.
Figure 2: Distribution of article-level annotation labels
Figure 3: Distribution of quantity-level annotation labels
@@ 4 Model
To model framing, we take advantage of the relations between the values of the frame components for a given article and implement a relational model using Probabilistic Soft Logic (PSL) Bach et al. (2017). Dependencies in PSL are expressed using weighted logical rules of the form: $w_{r}:P_{1}\wedge...\wedge P_{n-1}\rightarrow P_{n}$, where $w_{r}$ indicates the importance of the rule in the model, and can be learned from data. Predicates $P_{i}$ correspond to decisions and observations. Rules are then compiled into a Hinge-Loss Markov random field and weights are learned using maximum likelihood estimation. We consider the following rules:
Priors. We explicitly model prior probabilities for all article-level (type, economic conditions and economic direction) and quantity-level (type, macro-indicator, and sentiment) components. These priors are derived from supervised classifiers trained on our labeled dataset.
@@
$(r_{2})$ Consistency between the article type, economic conditions and direction. We enforce that if the article type is macro, then economic conditions and economic direction must be predicted and vice-versa: $\mathtt{Type(a,macro)}\Leftrightarrow\neg\mathtt{EconConditions(q,irrelevant)}\wedge\neg\mathtt{EconRating(q,irrelevant)}$
@@
$(r_{5})$ Dependency between neighboring quantity types. Our intuition is that there are common patterns in the types of consecutive quantities (e.g., sequential dependencies): $\mathtt{Precedes(q1,q2)}\wedge\mathtt{Type(q1,t1)}\Rightarrow\mathtt{Type(q2,t2)}$
@@
Pre-trained language models (LMs) are one of the most effective techniques for acquiring knowledge from unlabeled text data Devlin et al. (2019). Domain-adaptive pre-training (DAPT) further enriches LMs with in-domain data Gururangan et al. (2020). In both of these pre-training stages, learning is conducted by randomly masking words in a large dataset and training the the LM to predict these words. To take advantage of these strategies, we use pre-trained RoBERTA and perform DAPT with our large unlabeled dataset (all 199,066 economic articles collected). We use this LM to fine-tune the classifiers that are used as priors in PSL. This strategy has been repeatedly used to combine PSL with strong classifiers, with consistent success  Sridhar et al. (2015); Pacheco and Goldwasser (2021); Roy et al. (2021).
@@ 5 Experimental Evaluation
In this section, we outline our experimental settings and present results for our frame prediction model. We also include an example of the type of framing analysis that can be done using our model predictions over the complete dataset.